OpenStudy (anonymous):

Monte performs an experiment using 2 identical graduated cylinders with a radius of 1.75 cm. The volume of the liquid in the first graduated cylinder is 150 cm3. The volume of the liquid in the second graduated cylinder is 170 cm3. What is the difference in the height of the liquid in the two cylinders? Use 3.14 to approximate pi, and round your answer to one decimal place.

OpenStudy (igreen):

@cvliboi Use the formula for the volume of a sphere. \(V = \pi r^2 h\) Cylinder #1: \(V = \pi r^2 h\) \(150 = (3.14)(1.75^2)h\) Solve that for 'h' @cvliboi

OpenStudy (anonymous):

for 1.75 squared I got 3.0625?

OpenStudy (igreen):

Yep, now multiply 3.0625 * 3.14

OpenStudy (igreen):

Can you multiply those two numbers?

OpenStudy (igreen):

@cvliboi

OpenStudy (chunkymonkay):

@iGreen is the answer 9.61625?